OpenStudy (anonymous):

Use the Special Right Triangle to evaluate sin 60°, cos 60° and tan 60°. Your answers should be exact (not a decimal). A. sin 60 = 1/2, cos 60 = √3/2, tan 60 = √3/3 B. sin 60 = √2/2, cos 60 = √2/2, tan 60 = 1 C. sin 60 = 1, cos 60 = 0, tan 60 = undefined D. sin 60 = √3/2, cos 60 = 1/2, tan 60 = √3

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

for a 30-60-90 triangle the units are: 30 degrees: x 60 degrees: root 3 times x 90 degrees: 2x sin 60 = opposite/hypotenuse \[\sin 60 = \sqrt{3}/2\] cos 60 = adjacent/hypotenuse \[\cos 60 = x/2x = 1/2\] tan 60 = opposite/adjacent \[\tan 60 = \sqrt{3}/1 = \sqrt{3}\]